OpenStudy (anonymous):

Find the coordinates for the midpoint of the segment with endpoints given. (5, 6) and (8, 2)

OpenStudy (anonymous):

since m=n there fore, midpoint= \[\frac{ z _{2}+z _{1} }{ 2 } \] x coordinate =\[\frac{ 5+8 }{ 2}\] =13/2 y coordinate= \[\frac{ 6+2 }{ 2 }\] =4 midpoint= (13/2, 4)